You are the manager of a firm that sells its product in a competitive market at a price of $40. Your firm's cost function is C = 60 + 4Q2. Your firm's maximum profits are


a. 36


b. 40


c. 60


d. 80


1
Expert's answer
2021-11-16T11:36:21-0500

"C=60+4Q^2"

"P=" $"40"

Optimal price: $40.

Optimal output:

"MR=P=40"

"MC=8Q"

To get optimal output:

"MR=MC"

"40=8Q"

"Q=5"

Maximum profit:

"=TR-TC"

"=PQ-(60+4Q^2)"

"=(40\times5)-[60+4(5^2)]"

"=40."

The firm's maximum profits are: $40.